# Random Transverse Field Effects on Magnetic Noise in Spin Systems

### v1

@@ II.1 Ising chain and stochastic dynamics
where the $s_{i}=\pm 1$ are classical variables, and we have set the exchange coupling strength $J=1$ as the reference energy scale (we shall further assume $k_{B}=1$ for the Boltzmann constant). We endow this model with single-spin-flip Markovian stochastic dynamics, governed by Metropolis-Hastings (MH) acceptance probabilities [33]. We shall refer to this system as the kinetic Ising chain (KIC) [34, 35]. Our choice of stochastic discrete-time dynamics is equivalent to the one generated by the classical Master equation:
$\frac{\mathrm{d}}{\mathrm{d}t}P(\bm{\sigma})=\sum_{\bm{\sigma}^{\prime}}w(\bm{\sigma}^{\prime}|\bm{\sigma})P(\bm{\sigma}^{\prime})-w(\bm{\sigma}|\bm{\sigma}^{\prime})P(\bm{\sigma})\,,$ (2)
where $P(\bm{\sigma})$ denotes the statistical probability of the spin configuration $\bm{\sigma}$ and $w(\bm{\sigma}^{\prime}|\bm{\sigma})$ is the transition rate from $\bm{\sigma}^{\prime}$ to $\bm{\sigma}$, which is given by
$w(\bm{\sigma}|\bm{\sigma}^{\prime})=\left\{\begin{aligned} &\frac{1}{\tau_{0}}\min\left(1,e^{-\beta(E_{\bm{\sigma}^{\prime}}-E_{\bm{\sigma}})}\right),&\bm{\sigma}^{\prime}=S_{i}^{x}\bm{\sigma}\,,\\ &0,&\text{otherwise.}\end{aligned}\right.$ (3)
where, in this classical model, $S_{i}^{x}$ represents the action of flipping the spin at the $i$th site. The characteristic time scale $\tau_{0}$ is the microscopic single-spin-flip time, usually set to $1$ in MH dynamics.
@@ II.2 Random transverse field
In order to develop a dynamical version of the RTFIC, we need to generalize the classical MH dynamics to account for the presence of the quantum clusters. This is done by coupling the system described by Eq. (4) to an external heat bath at a given temperature and inducing dissipative dynamics. Correspondingly, the reduced density matrix of the system, $\rho$, follows the quantum master equation [37, 38]
@@
where the $|a\rangle$ and $|b\rangle$ are eigenstates of the system Hamiltonian and $\delta_{\omega^{\prime},\omega}$ is a Kronecker delta. The heat bath attempts to flip a spin at site $i$ by applying the Pauli operator $S_{i}^{x}$. To maintain detailed balance, Eq. (8) decomposes this operator in the basis of system eigenstates and selects the part that connects two states with energy difference $\omega$, which is then scaled by a factor of the system-bath coupling strength $\gamma_{i}(\omega)$ for this transition mode. Detailed balance is obeyed for $\gamma_{i}(\omega)=e^{-\beta\omega}\gamma_{i}(-\omega)$.
Figure 1: Magnetic noise spectrum of the RTFIC: (a) PSDs of the RTFIC extracted using Welch’s method on MC simulations, with $r=0.2$ and varying values of $h$ and $T$. PSDs for the KIC are also shown for reference. Inset: $\omega^{2}\,\mathrm{PSD}(\omega)$ as a function of $\omega$, with $r=0.2$ and varying values of $h$ and $T=0.3$, focusing on the intermediate region connecting the two inverse-square scaling regions. (b) Values for the parameter $A$, obtained by fitting the two regions of the PSD bounded by the dashed vertical lines in (a) to the form $\mathrm{PSD}(\omega)=A\omega^{-2}$, at $r=0.1$ (blue) or $0.2$ (red). While the low-frequency (dash-dot) and high-frequency (solid) results overlap for the KIC, they differ for the RTFIC, with the discrepancy becoming more pronounced with increasing $h,\,r$ and decreasing $T$. In both figures, results for different temperatures are shifted vertically for clarity.
@@
where $|b\rangle$, $|c\rangle$ are states drawn from the Hamiltonian eigenbasis, and we have introduced the short-hand notation $M_{i,abc}\equiv\langle a|S_{i}^{x}|b\rangle\langle b|S_{i}^{x}|c\rangle\delta_{E_{a},E_{c}}$. For the KIC, we select the eigenbasis $\{|\alpha\rangle\}$ to be identical to the spin-configuration basis $\{\bm{\sigma}\}$, under which we have $M_{i,abc}=\delta_{a,c}\delta_{b,S_{i}^{x}a}$. In this case, Eq. (9) reduces to the classical Master equation, Eq. (2), by identifying $\langle\bm{\sigma}|\rho|\bm{\sigma}\rangle=P(\bm{\sigma})$ under the selection $\gamma_{i}(\omega)=\min(1,\mathrm{e}^{-\beta\omega})$ where the single-spin-flip time $\tau_{0}$ has been set to $1$. For the RTFIC, we select $\{|\alpha\rangle\}$ to be the set of the product states introduced above. Now, the question arises whether one still has $M_{i,abc}\propto\delta_{a,c}$ so that the RHS of Eq. (9) involves only diagonal elements of the density matrix. In App. B we show that, although there exist some rare cases where the above proportionality is violated, they have low statistical weight at low temperature and can be safely neglected. Using the approximation $M_{i,abc}=\left|\langle a|S_{i}^{x}|b\rangle\right|^{2}\delta_{a,c}$, we arrive at the following Master equation for the RTFIC:
@@ III.1 Magnetic noise spectrum
In the limit of isolated kink or anti-kink motion in the classical system ($r=0$), the mapping to a random walk allows to derive the known scaling result $C(\omega)\propto\omega^{-2}$ [41]. This provides a starting point to study the behavior of the RTFIC when the randomly walking domain walls are affected by the presence of quantum clusters.
The results for the PSD, obtained using Welch’s method [42] on MC simulations with different values of $r$, $h$ and $T$, are presented in Fig. 1a. (For details of the parameters used to run the MC simulations, see App. F.) A deviation from the $\omega^{-2}$ behavior of the KIC at low frequency is observed around $\omega\sim 10^{-1}$ (in units of $2\pi/\text{MC time}$), when the PSD of the RTFIC exhibits a plateau-like upturn which terminates at $\omega\sim 10^{1}$, followed by another inverse square scaling regime, $\mathrm{PSD}(\omega)\propto\omega^{-2}$, with different proportionality factor depending on the value of $h$.
We fit the function $\mathrm{PSD}(\omega)=A\omega^{-2}$ to the low- and high-frequency part of the curve and extract the proportionality factors $A_{0}=\lim_{\omega\to 0}\omega^{2}\,\mathrm{PSD}(\omega)$ and $A_{\infty}=\lim_{\omega\to\infty}\omega^{2}\,\mathrm{PSD}(\omega)$ as functions of the parameters $r$, $h$ and $T$. As seen in Fig. 1b, the effect of the quantum clusters is most prominent in the high-frequency scaling regime of the PSD, increasing with $h,\,r$ and decreasing with $T$. Note that the results for the PSD or $A$ do not converge to the classical result as $h\to 0$ for reasons that we shall discuss below.
@@ III.2 Quasiparticle random walk dynamics
Figure 2: Quasiparticle random-walk dynamics: (a) Illustration of the random walk model. We associate each pair of nearest-neighbor classical spins and each cluster with a single site on the RW lattice. An excitation can manifest itself either classically (as anti-aligned spin pairs) or as a cluster with anti-aligned boundary spins. These excitations move by flipping classical spins in the system. Labelings of the quantum clusters follow the convention given in App. A. (b) Results for the MSD, $M(t)$, with $T=0.3,\,r=0.2$. One sees the emergence of two distinct diffusive regimes for $t\to 0$ and $t\to\infty$ as $h$ increases, deviating (in both cases) from the KIC, where $M(t)$ exhibits the same linear scaling throughout. We also plot the fits to the data with $h=2.4$ according to Eq. (17). Inset: the diffusion constants $D_{0}$ (empty markers) and $D_{\infty}$ (filled markers), plotted as a function of $h$ and $T$, for $r=0.2$. They both decrease as $h$ increases, yet their difference increases with $h$.
@@
In our case, $X(t)$ is identified as the magnetization of the system. The modifications of the PSD seen in the previous section due to the quantum clusters can then be attributed to the interplay between the clusters and the mobile RW excitations, which alters $M(t)$.
@@
Each site of the RW lattice corresponds to a possible position of a single excitation, which can be mapped to the magnetization of the system injectively (see App. C). In the KIC, excitations manifest themselves classically as domain walls, i.e., anti-aligned nearest-neighbor classical spin pairs. On the RW lattice, this corresponds to a ‘classical’ site in the middle of the corresponding bond. When quantum clusters are present, excitations may also manifest themselves as clusters with anti-aligned boundary spins, see Fig. 2a. The cases $h=0$ and finite $h$ are drastically different for the clusters: at $h=0$, the ground state of such a cluster with anti-aligned boundary spins is $n+1$-fold degenerate, corresponding to the $n+1$ possible positions of a single domain wall; any infinitesimal but nonzero $h$ lifts this degeneracy, creating quantum superpositions that necessitate the cluster to be treated as a single site. Such modification of the RW lattice for arbitrary $h>0$ explains the deviation between the KIC and the RTFIC for $h\to 0$ mentioned above.
Motion on the sites of the RW lattice takes place by flipping spins. In between classical sites, this process is the same in the RTFIC as in the KIC: domain walls move by one lattice spacing when one of their spins is flipped, inducing a magnetization change $\Delta m=2$. On the contrary, moving into, out of and in between quantum clusters is accomplished by flipping one of the boundary spins. Such moves can involve several sites of the original lattice, and non-trivial changes in magnetization; according to our mapping, however, these processes still correspond to nearest-neighbor motion on the RW lattice. This subtlety will require careful modeling in the following.
@@
where $P_{i}$ denotes the probability of finding the walker at site $i$ of the RW lattice, and $W_{ij}$ is the transition rate from site $i$ to site $j$. The latter is related to the MH probability, Eq. (12), of the spin flip that induces the hopping from site $i$ to site $j$. Such probability is $1$ when both sites are classical, giving $W_{ij}=1$; in all other cases, $W_{ij}$ take a value determined by the complex energy landscapes of the relevant quantum clusters. These values are explicitly evaluated in App. C.
As explained above, not all nearest-neighbor hops on the RW lattice result in the same change in magnetization of the system. It is therefore necessary to complement the Master equation, Eq. (16), with a way to account for the size and energy level structure of the quantum clusters. This is equivalent to a heterogeneous RW lattice with varying step lengths across the quantum clusters. In App. C, we present details of the evaluation of the magnetization changes and corresponding step lengths.
@@
By fitting the asymptotic linear behavior, $\lim_{t\to 0}M(t)=D_{0}t$ and $\lim_{t\to\infty}M(t)=D_{\infty}t$, we obtain the diffusion constants $D_{0}$ and $D_{\infty}$. For the KIC, where the excitations perform a conventional random walk, we have $M(t)=Dt$ with $D=8$ at all times (where a factor of $2$ comes from coordination number, and another factor of $42^{2}=$ 4 comes from the unit step length $\Delta m=2$ induced by flipping a classical spin and the square scaling relationship between the diffusion constants and the unit bond length [45]). For the RTFIC, results for $D_{0}$ and $D_{\infty}$ for varying $h$ and $T$ are shown in the inset of Fig. 2b, where we see the most prominent deviations between $D_{0}$ and $D_{\infty}$ at large $h$ and small $T$. Note that, as mentioned before, the results for $h=0^{+}$ for the RTFIC are not continuously connected with the behaviour of the KIC, due to the degeneracy-lifting effect of the transverse fields.
Figure 3: Results for $D_{0}/D_{\infty}$ from the random walk model (round points, solid lines) and $A_{\infty}/A_{0}$ extracted from computing the PSD of the random walk (rhombic points, dashed lines) and the full-scale spin simulations (triangular points, dotted lines) presented in Fig. 1. The horizontal dotted line represents the value of $A_{\infty}/A_{0}=D_{0}/D_{\infty}=1$ for the case of KIC. Consistency between the three groups of data validates the single-excitation picture and the approximation of the form Eq. (17) used in the main text. Inset: the same data plotted against $h/T$ rather than $h$; they scale with $h/T$ at small values, while they deviate at large values.
@@ III.3 Discussion
The above investigations highlight the usefulness of two key quantities — $A_{\infty}/A_{0}$, derived from the magnetic noise of the system, and $D_{0}/D_{\infty}$, obtained from studying the single-excitation dynamics. Both take the trivial value of $1$ in the KIC and can serve to quantify deviations from diffusive behaviors. A full analytical solution for the diffusion constants $D_{0}$ and $D_{\infty}$ is provided in Appendix C.
@@ IV.1 Model
We consider the RTFIC, Eq. (4), under a sinusoidally oscillating longitudinal driving field
@@
In the driven model, the system responds to the oscillating field $B(t)$ with a time-dependent magnetization of the form $m(t)=m_{0}\sin(\omega t+\phi)$, where $\phi$ is the phase shift. We are interested in the magnetic susceptibility of the system, $\chi(\omega)=m_{0}\mathrm{e}^{i\phi}/B_{0}$ and its real and imaginary components, denoted by $\chi^{\prime}(\omega)$ and $\chi^{\prime\prime}(\omega)$ respectively. These quantities are closely related to the PSD and the MSD that we studied in the previous section via the fluctuation-dissipation theorem. In particular, for single excitations, we have
@@
We start with the KIC where, in the single-excitation limit, $M(t)$ is linear in time. Performing the appropriate integration in Eq. (20) yields $\mathrm{PSD}(\omega)\propto\omega^{-2}$, and therefore an $\omega$-independent $\omega\chi^{\prime\prime}(\omega)$. Furthermore, Eq. (21) yields $\chi^{\prime}(\omega)=0$, giving $\phi=\pi/2$.

### v2

@@ II.1 Ising chain and stochastic dynamics
where the $s_{i}=\pm 1$ are classical variables, and we have set the exchange coupling strength $J=1$ as the reference energy scale (we shall further assume $k_{B}=1$ for the Boltzmann constant). We endow this model with single-spin-flip Markovian stochastic dynamics, governed by Metropolis-Hastings acceptance probabilities [33]. Under such dynamics, each spin attempts to flip on a characteristic time scale which is the microscopic single-spin-flip time, conventionally set to $1$ in Metropolis-Hastings dynamics. The spin-flip events are accepted with probability $P(\bm{\sigma}\to\bm{\sigma}^{\prime})=\min\left(1,\mathrm{e}^{-\beta\Delta E}\right)$, where $\beta=1/T$ is the inverse temperature and $\Delta E$ denotes the energy difference between the two spin configurations $\bm{\sigma}$ and $\bm{\sigma}^{\prime}$. We shall refer to this system as the kinetic Ising chain [34, 35]. Our choice of stochastic discrete-time dynamics is equivalent to the one generated by the classical Master equation.
@@ II.2 Random transverse field
In order to develop a dynamical version of the RTFIC, we need to generalize the classical Metropolis-Hastings dynamics to account for the presence of the quantum clusters. This is done by coupling the system described by Eq. (2) to an external heat bath at a given temperature and inducing dissipative dynamics. Correspondingly, the reduced density matrix of the system, $\rho$, follows the quantum master equation [37, 38]
@@
[…]
Figure 1: Magnetic noise spectrum of the RTFIC: (a) PSDs of the RTFIC extracted using Welch’s method on MC simulations, with $r=0.2$ and varying values of $h$ and $T$. PSDs for the classical kinetic Ising model are also shown for reference. Inset: $\omega^{2}\,\mathrm{PSD}(\omega)$ as a function of $\omega$, with $r=0.2$ and varying values of $h$ and $T=0.3$, focusing on the intermediate region connecting the two inverse-square scaling regions. (b) Estimates for the values of the parameters $A_{\infty}$ and $A_{0}$, as defined in the main text, obtained by fitting the two regions of the PSD bounded by the dashed vertical lines in (a) to the form $\mathrm{PSD}(\omega)=A\omega^{-2}$, at $r=0.1$ (blue) or $0.2$ (red). While the low-frequency (dotted) and high-frequency (solid) results shall overlap for the classical model, they differ for the RTFIC, with the discrepancy becoming more pronounced with increasing $h,\,r$ and decreasing $T$. In both figures, results for different temperatures are shifted vertically for clarity.
@@
where $|b\rangle$, $|c\rangle$ are states drawn from the Hamiltonian eigenbasis, and we have introduced the short-hand notation $M_{i,abc}\equiv\langle a|S_{i}^{x}|b\rangle\langle b|S_{i}^{x}|c\rangle\delta_{E_{a},E_{c}}$. For the classical kinetic Ising chain, we select the eigenbasis $\{|\alpha\rangle\}$ to be identical to the spin-configuration basis $\{\bm{\sigma}\}$, under which we have $M_{i,abc}=\delta_{a,c}\delta_{b,S_{i}^{x}a}$. In this case, Eq. (7) reduces to the classical Master equation, by identifying $\langle\bm{\sigma}|\rho|\bm{\sigma}\rangle=P(\bm{\sigma})$ as the statistical probability of the spin configuration $\bm{\sigma}$ and under the selection $\gamma_{i}(\omega)=\min(1,\mathrm{e}^{-\beta\omega})$ where the single-spin-flip time $\tau_{0}$ has been set to $1$. For the RTFIC, we select $\{|\alpha\rangle\}$ to be the set of the product states introduced above. Now, the question arises whether one still has $M_{i,abc}\propto\delta_{a,c}$ so that the RHS of Eq. (7) involves only diagonal elements of the density matrix. In App. B we show that, although there exist some rare cases where the above proportionality is violated, they have low statistical weight at low temperature and can be safely neglected. Using the approximation $M_{i,abc}=\left|\langle a|S_{i}^{x}|b\rangle\right|^{2}\delta_{a,c}$, we arrive at the following Master equation for the RTFIC:
@@ III.1 Magnetic noise spectrum
In the limit of isolated kink or anti-kink motion in the classical system ($r=0$), the mapping to a random walk allows to derive the known scaling result ${\mathrm{PSD}}(\omega)\propto\omega^{-2}$ [43]. This provides a starting point to study the behavior of the RTFIC when the randomly walking domain walls are affected by the presence of quantum clusters.
The results for the PSD, obtained using Welch’s method [44] on MC simulations with different values of $r$, $h$ and $T$, are presented in Fig. 1a. (For details of the parameters used to run the MC simulations, see App. F.) A deviation from the classical $\omega^{-2}$ behavior at low frequency is observed around $\omega\sim 10^{-1}$ (in units of $2\pi/\text{MC time}$), when the PSD of the RTFIC exhibits a plateau-like upturn which terminates at $\omega\sim 10^{1}$, followed by another inverse square scaling regime, $\mathrm{PSD}(\omega)\propto\omega^{-2}$, with different proportionality factor depending on the value of $h$.
We fit the function $\mathrm{PSD}(\omega)=A\omega^{-2}$ to the low- and high-frequency part of the curve (see the dashed vertical lines in Fig. 1), and use the values obtained for $A$ as estimates of the proportionality factors $A_{0}=\lim_{\omega\to 0}\omega^{2}\,\mathrm{PSD}(\omega)$ and $A_{\infty}=\lim_{\omega\to\infty}\omega^{2}\,\mathrm{PSD}(\omega)$. We plot these values as functions of the parameters $r$, $h$ and $T$. As seen in Fig. 1b, while these two values agree in the classical scenario, they differ in the RTFIC; the effect of the quantum clusters is most prominent in the high-frequency regime of the PSD, as $A_{\infty}$ has a more pronounced dependence on $h$ and $r$ comparing with $A_{0}$. The difference between these two quantities increases with $h,\,r$ and decreases with $T$. In addition, we note that the results obtained in the limit $h\to 0$ (see App. F) give values of $A$ that do not converge to the classical result. We shall discuss the reason for this discrepancy later in the manuscript.
@@ III.2 Quasiparticle random walk dynamics
Figure 2: Quasiparticle random walk dynamics: (a) Illustration of the random walk model. We associate each pair of nearest-neighbor classical spins and each cluster with a single site on the random walk lattice. An excitation can manifest itself either classically (as anti-aligned spin pairs) or as a cluster with anti-aligned boundary spins. These excitations move by flipping classical spins in the system. Labeling of the quantum clusters follow the convention given in App. A. (b) Results for the MSD, $M(t)$, with $T=0.3,\,r=0.2$. One sees the emergence of two distinct diffusive regimes for $t\to 0$ and $t\to\infty$ as $h$ increases, deviating (in both cases) from the red classical scenario, where $M(t)$ exhibits the same linear scaling throughout. We also plot the fits to the data with $h=2.4$ according to Eq. (15). Inset: estimates of the diffusion constants $D_{0}$ (empty markers) and $D_{\infty}$ (filled markers), plotted as a function of $h$ and $T$, for $r=0.2$. They both decrease as $h$ increases, yet their difference increases with $h$.
@@
In our case, $X(t)$ is identified as the magnetization of the system. The modifications of the PSD seen in the previous section due to the quantum clusters can then be attributed to the interplay between the clusters and the mobile random walk excitations, which alters $M(t)$.
@@
Each site of the random walk lattice corresponds to a possible position of a single excitation, which can be mapped to the magnetization of the system injectively (see App. C). In the classical scenario, excitations manifest themselves classically as domain walls, i.e., anti-aligned nearest-neighbor classical spin pairs. On the random walk lattice, this corresponds to a ‘classical’ site in the middle of the corresponding bond. When quantum clusters are present, excitations may also manifest themselves as clusters with anti-aligned boundary spins, see Fig. 2a. The cases $h=0$ and finite $h$ are drastically different for the clusters: at $h=0$, the ground state of such a cluster with anti-aligned boundary spins is $n+1$-fold degenerate, corresponding to the $n+1$ possible positions of a single domain wall; any infinitesimal but nonzero $h$ lifts this degeneracy, creating quantum superpositions that necessitate the cluster to be treated as a single site. Such modification of the random walk lattice for arbitrary $h>0$ explains the deviation between the RTFIC at $h\to 0$ to the classical scenario as mentioned above.
Motion on the sites of the random walk lattice takes place by flipping spins. In between classical sites, domain walls move by one lattice spacing when one of their spins is flipped, inducing a magnetization change $\Delta m=2$. On the contrary, moving into, out of and in between quantum clusters is accomplished by flipping one of the boundary spins. Such moves can involve several sites of the original lattice, and non-trivial changes in magnetization; according to our mapping, however, these processes still correspond to nearest-neighbor motion on the random walk lattice. This subtlety will require careful modeling in the following.
@@
where $P_{i}$ denotes the probability of finding the walker at site $i$ of the random walk lattice, and $W_{ij}$ is the transition rate from site $i$ to site $j$. The latter is related to the Metropolis-Hastings probability, Eq. (10), of the spin flip that induces the hopping from site $i$ to site $j$. Such probability is $1$ when both sites are classical, giving $W_{ij}=1$; in all other cases, $W_{ij}$ take a value determined by the complex energy landscapes of the relevant quantum clusters. These values are explicitly evaluated in App. C.
As explained above, not all nearest-neighbor hops on the random walk lattice result in the same change in magnetization of the system. It is therefore necessary to complement the Master equation, Eq. (14), with a way to account for the size and energy level structure of the quantum clusters. This is equivalent to a heterogeneous random walk lattice with varying step lengths across the quantum clusters. In App. C, we present details of the evaluation of the magnetization changes and corresponding step lengths.
@@
By fitting the asymptotic linear behavior, $\lim_{t\to 0}M(t)=D_{0}t$ and $\lim_{t\to\infty}M(t)=D_{\infty}t$, we obtain the diffusion constants $D_{0}$ and $D_{\infty}$. In the classical scenario, where the excitations perform a conventional random walk, we have $M(t)=Dt$ with $D=8$ at all times (where a factor of $2$ comes from coordination number, and another factor of $42^{2}=$ 4 comes from the unit step length $\Delta m=2$ induced by flipping a classical spin and the square scaling relationship between the diffusion constants and the unit bond length [47]). For the RTFIC, results for $D_{0}$ and $D_{\infty}$ for varying $h$ and $T$ are shown in the inset of Fig. 2b, where we see the most prominent deviations between $D_{0}$ and $D_{\infty}$ at large $h$ and small $T$. Note that, as mentioned before, the results for $h=0^{+}$ for the RTFIC are not continuously connected with the classical behavior, due to the degeneracy-lifting effect of the transverse fields.
Figure 3: Results for $D_{0}/D_{\infty}$ from the random walk model (round points, solid lines) and $A_{\infty}/A_{0}$ extracted from computing the PSD of the random walk (rhombic points, dashed lines) and the full-scale spin simulations (triangular points, dotted lines) presented in Fig. 1. The horizontal dotted line represents the value of $A_{\infty}/A_{0}=D_{0}/D_{\infty}=1$ for the classical case. Consistency between the three groups of data validates the single-excitation picture and the approximation of the form Eq. (15) used in the main text. Inset: the same data plotted against $h/T$ rather than $h$; they scale with $h/T$ at small values, while they deviate at large values.
@@ III.3 Discussion
The above investigations highlight the usefulness of two key quantities — $A_{\infty}/A_{0}$, derived from the magnetic noise of the system, and $D_{0}/D_{\infty}$, obtained from studying the single-excitation dynamics. Both take the trivial value of $1$ in the classical scenario and can serve to quantify deviations from diffusive behaviors. A full analytical solution for the diffusion constants $D_{0}$ and $D_{\infty}$ is provided in Appendix C.
@@ IV.1 Model
We consider the RTFIC, Eq. (2), under a sinusoidally oscillating longitudinal driving field [55]
@@
In the driven model, the system responds to the oscillating field $B(t)$ with a time-dependent magnetization of the form $m(t)=m_{0}\sin(\omega t+\phi)$, where $\phi$ is the phase shift. We are interested in the magnetic susceptibility of the system, $\chi(\omega)=m_{0}\mathrm{e}^{i\phi}/B_{0}$ and its real and imaginary components, denoted by $\chi^{\prime}(\omega)$ and $\chi^{\prime\prime}(\omega)$ respectively. These quantities are closely related via the fluctuation-dissipation theorem to the PSD of the system and, at low temperatures, to the MSD of single excitations that we studied in the previous section. In particular, we have
@@
We start with the classical version of Eq. (17), where, for a single mobile excitation, $M(t)$ is linear in time. Performing the appropriate integration in Eq. (18) yields $\mathrm{PSD}(\omega)\propto\omega^{-2}$, and therefore an $\omega$-independent $\omega\chi^{\prime\prime}(\omega)$. Furthermore, Eq. (19) yields $\chi^{\prime}(\omega)=0$, giving $\phi=\pi/2$.